Two pulleys with masses m1 = 1 (kg) and m2 = 2 (kg) fixed
coaxially to each other are fixed on the wall so that they can rotate around point O (bigger pulley has 2r and the small one on the inside of the big one has radius r). Mass m3 = 3 (kg) is connected to the rope wound on a reel with radius r = 1 (m). Initially a constant force of
1
Expert's answer
2020-12-23T07:35:31-0500

Moment of inertia of system can be written as

"I=\frac{1}{2}m_1(2r) ^2+\frac{1}{2}m_2r^2"

By putting value

"I=\frac{1}{2}(1)(2r) ^2+\frac{1}{2}(2)r^2\\I=3r^2"